# Optimal Threshold for Fracton Codes and Nearly Saturated Code Capacity in Three Dimensions

### v1

@@ VI Summary
In this work, we investigated the fault-tolerance threshold of self-dual fracton codes. By mapping the error correction problem onto a statistical mechanical model and performing large-scale numerical simulations, we computed the optimal thresholds of the checkerboard code against stochastic Pauli noise. Our numerical analysis identifies a code capacity of $p_{th}\simeq 0.108(2)$ for the checkerboard code. This value is notable not only for its magnitude—outperforming other $3D$ codes—but because it represents the first instance of a $3D$ code that nearly saturates the theoretical limit of $p_{th}\approx 0.11$ allowed by the quantum Gilbert-Varshamov bound [Varshamov57, Calderbank96].
@@
Moreover, our results further corroborate the generalized entropy duality in Eq.(16), confirming its applicability beyond standard topological codes. The remarkable effectiveness of this duality may be attributed to the self-averaging nature of both the ordered and disordered paramagnetic phases and the lack of spin glass phase along the Nishimori line. This ensures that the self-consistent replica analysis yields reasonable descriptions. Our validation carries significant practical importance alongside its theoretical implications. Exact threshold computations are notoriously resource-intensive and generally lack efficient algorithms. For $3D$ codes in particular, determining their thresholds can consume millions of CPU hours. The generalized duality can save immense computational resources. Leveraging this predictive power, we conjecture that the code capacity of Haah’s code, which is even more computationally demanding to simulate, also approaches the theoretical limit of $p_{th}\approx 0.11$. Similarly, one can also expect the bivariate bicycle codes [Bravyi24], which are a rich class quantum low-density parity-check codes characterized by symmetry-enriched topological order [Chen25], also possess such an optimal code capacity.
@@
Looking forward, while this work focused on code capacity, a critical next step is to extend the analysis to include measurement errors and circuit-level noise. Developing effective strategies to treat these more general noise models beyond the surface code remains an exciting open question. Furthermore, from a theoretical perspective, it is of great interest to investigate the precise nature of the correctable-to-uncorrectable phase transition across the threshold, which may reveal novel critical behaviors unique to fracton states of matter.
@@
Acknowledgments. G.C., K.L., and L.P. acknowledge support from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC-2111 – 390814868. MAMD acknowledges support from Spanish MICIN grant PID2021-122547NB-I00 and the “MADQuantumCM” project funded by Comunidad de Madrid (Programa de acciones complementarias) and by the Ministry for Digital Transformation and of Civil Service of the Spanish Government through the QUANTUM ENIA project call –Quantum Spain project, and by the European Union through the Recovery, Transformation and Resilience Plan Next Generation EU within the framework of the Digital Spain 2026 Agenda, the CAM Programa TEC-2024/COM-84 QUITEMAD-CM. H.S. acknowledges support from the National Natural Science Foundation of China (Grants No. 12522502, No. 12474145, and No. 12447101). The project/research is part of the Munich Quantum Valley, which is supported by the Bavarian state government with funds from the Hightech Agenda Bayern Plus.

### v2

@@ VI Summary
In this work, we investigated the fault-tolerance threshold of self-dual fracton codes. By mapping the error correction problem onto a statistical mechanical model and performing large-scale numerical simulations, we computed the optimal thresholds of the checkerboard code against stochastic Pauli noise. Our numerical analysis identifies a code capacity threshold of $p_{th}\simeq 0.107(3)$ for the checkerboard code. This value is notable not only for its magnitude—outperforming other $3D$ codes—but because it represents the first instance of a $3D$ code that nearly saturates the theoretical limit of $p_{th}\approx 0.11$ allowed for topological stabilizer codes [20].
@@
Moreover, our results further corroborate the generalized entropy duality in Eq.(16), confirming its applicability beyond standard topological codes. The remarkable effectiveness of this duality may be attributed to the self-averaging nature of both the ordered and disordered paramagnetic phases and the lack of spin glass phase along the Nishimori line. This ensures that the self-consistent replica analysis yields reasonable descriptions. Our validation carries significant practical importance alongside its theoretical implications. Exact threshold computations are notoriously resource-intensive and generally lack efficient algorithms. For $3D$ codes in particular, determining their thresholds can consume millions of CPU hours. The generalized duality can save immense computational resources. Leveraging this predictive power, we conjecture that the code capacity threshold of Haah’s code, which is even more computationally demanding to simulate, also approaches the theoretical limit of $p_{th}\approx 0.11$. Similarly, one can also expect the bivariate bicycle codes [8], which are a rich class quantum low-density parity-check codes characterized by symmetry-enriched topological order [18], also possess such an optimal error threshold.
@@
Looking forward, while this work focused on code capacity threshold, a critical next step is to extend the analysis to include measurement errors and circuit-level noise. Developing effective strategies to treat these more general noise models beyond the surface code remains an exciting open question. Furthermore, from a theoretical perspective, it is of great interest to investigate the precise nature of the correctable-to-uncorrectable phase transition across the threshold, which may reveal novel critical behaviors unique to fracton states of matter.
@@
Acknowledgments. G.C., K.L., and L.P. acknowledge support from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC-2111 – 390814868. MAMD acknowledges support from Spanish MICIN grant PID2021-122547NB-I00 and the “MADQuantumCM” project funded by Comunidad de Madrid (Programa de acciones complementarias) and by the Ministry for Digital Transformation and of Civil Service of the Spanish Government through the QUANTUM ENIA project call –Quantum Spain project, and by the European Union through the Recovery, Transformation and Resilience Plan Next Generation EU within the framework of the Digital Spain 2026 Agenda, the CAM Programa TEC-2024/COM-84 QUITEMAD-CM. H.S. acknowledges support from the National Natural Science Foundation of China (Grants No. 12522502, No. 12474145, and No. 12447101). L.P. acknowledges financial support from ANR-23-CE30-0018. The project/research is part of the Munich Quantum Valley, which is supported by the Bavarian state government with funds from the Hightech Agenda Bayern Plus.
Data availability. The Monte Carlo simulation data supporting the findings of this study are available in the repository indicated in Ref. [16].